OpenStudy (anonymous):

The formula for degrees Celsius is: C = (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is.

OpenStudy (anonymous):

that s not correct formula

OpenStudy (anonymous):

it should be C = (5/9) ( F - 32)

OpenStudy (anonymous):

BUT AS AN EXERCISE IN REARRANGING FORMULAE we can use it C = (F − 32) C = F - 32 add 32 to both sides C + 32 = F

OpenStudy (anonymous):

using this formula for 8 degrees centigrade Ff = 32 + 8 = 40

OpenStudy (anonymous):

yes 5/9 is suppose to be there

OpenStudy (anonymous):

sorry gotta go immediately

OpenStudy (anonymous):

C = (5/9) ( F - 32) 9 C = 5 F - 160 5F = 9C +160 F=( 9/5) C + 32